# Quantifying the generation of negatively charged boron vacancies in He-ion irradiated hexagonal boron nitride

### v1

@@ Abstract
Hexagonal boron nitride (hBN) hosts luminescent defects possessing spin qualities compatible with quantum sensing protocols at room temperature. Vacancies, in particular, are readily obtained via exposure to high-energy ion beams. While the defect creation mechanism via such irradiation is well understood, the occurrence rate of optically active negatively charged vacancies (V–B) is an open question. In this work, we exploit focused helium ions to systematically generate optically active vacancy defects in hBN flakes at varying density. By comparing the density-dependent spin splitting measured by magnetic resonance to calculations based on a microscopic charge model, in which we introduce a correction term due to a constant background charge, we are able to quantify the number of V–B defects generated by the ion irradiation. We find that only a small fraction (0.2%) of all vacancies is in the optically active, negatively charged state. Our results provide a protocol for measuring the generation efficiency of V–B, which is necessary for understanding and optimizing luminescent centers in hBN.
@@ I Introduction
Luminescent centers in hBN emerged as bright and stable room temperature quantum emitters [1, 2], rivaling established platforms in quantum communication and sensing applications [3, 4, 5]. Emitters in hBN exhibit robust photoluminescence (PL) across a broad spectral range from the near-infrared [6] through the visible [1, 7] up to the blue [8, 9, 10, 11, 12] and UV [13] regions, as well as magnetic response both as ensemble spin-1 systems [6] and as single spin-1 [14, 15] or spin-1/2 [16] emitters. Among the most studied types of luminescent centers in hBN is the negatively charged boron vacancy (V–B), which consists of an electron localized at the position of the missing atom in the hexagonal lattice. These defects are typically addressed in PL ensemble measurements, where they exhibit a very broad spectrum centered around 800 nm [6, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Importantly, V–B defects feature a spin-triplet ground state with a zero-field splitting of around 3.5 GHz [6], a corresponding triplet excited state with a splitting of around 2.1 GHz [25, 31, 32, 33], and a metastable state acting as a non-radiative channel with a very short lifetime [29, 30]. The spin state can be optically read-out and coherently manipulated through optically detected magnetic resonance (ODMR) measurement schemes [6, 28, 34, 35, 30]. The microwave resonances are sensitive to external magnetic fields [6, 25], which can be applied for quantum sensing schemes [36, 37, 38, 18], similar to nitrogen vacancy-centers in diamond [39, 40]. Because of their potential for applications in quantum technologies, the optical and magnetic properties of V–B defects have been a subject of intensive research, including their optical response at room [6, 25, 41, 28, 30] and cryogenic [20, 41] temperatures, as well as ODMR [6, 20, 31, 33], electron paramagnetic resonance (EPR) [6, 42], and spin-coherence measurements [4, 43, 44, 45]. Furthermore, several studies discuss various sample preparation and optimization protocols, such as irradiation sputtering [22], ion beam fluence and energy tuning [21, 41], high-temperature annealing [46, 23, 24], water exposure [22], as well as the integration of defects into photonic circuits, such as nanocavities [47, 48], plasmonic- [19] and metasurfaces [49].
@@
For photonic integration reliable, quantitative and deterministic methods for defect generation are desirable. In this context, different methods were successfully applied to create V–B in hBN, which include focused ion beam irradiation using H [17], He [46, 28, 41, 19, 23, 21, 50], N [30, 21, 20], Ga [6, 31], C [21, 25], Xe [20] and Ar [21, 20], neutron irradiation [27, 6, 38, 51, 52], femtosecond laser writing [53], and electron beam irradiation [42, 54]. In particular, several works discussed the usage of a focused He-ion beam source to effectively generate V–B defects in hBN [28, 41, 19, 23, 21, 50]. Focused He-ion beam exposure offers a nominal lateral defect positioning accuracy below 10 nm in very thin films [55], while preserving the local crystal structure of hBN at adequate doses. All such nanofabrication measures are crucial factors for an efficient integration of such luminescent centers in future photonic [56] and quantum imaging devices [18]. In this context, several works [41, 28, 23] already provided initial insights onto the He-ion generation process and the optical response of the generated ensemble emitters, but a systematic understanding, in particular a reliable quantification of defect density and its impact on the magnetic and optical properties of the generated emitters, is still lacking.
@@
In the present work, we use a helium ion microscope (HIM) to irradiate hBN flakes, deposited on thick graphite flakes, to generate V–B defects (Figure 1 a). The thick graphite acts as both a mechanical and optical protection layer, reducing the impact of back-scattered ions from the SiO2 substrate and suppressing unwanted background luminescence from the irradiation damage in the substrate. We systematically perform spatially- and time-resolved PL as well as ODMR measurements on defect ensembles for varying irradiation fluences, confirming the optical and magnetic activity of the generated emitters. We show that the focused He-ion beam technique can site-selectively produce luminescent and background-free emitters, in good agreement with the existing literature [6, 28, 41, 19, 23, 21, 50]. Furthermore, we apply a microscopic charge model [44, 52] to describe the defect density-dependent shift of the ODMR resonance frequencies. By introducing a background charge correction, we can quantify from the experimental data the density of the charged defects generated by the ion exposure. By comparison to molecular dynamics (MD) simulations, which estimate the number of total vacancies generated, we are able to quantify an overall efficiency of the charged defect creation process based on the focused He-ion beam irradiation. Our work provides a systematic study of defect production efficiency by comparing different irradiation fluences on homogeneous and uniform samples, offering a benchmark for future process optimization and device implementation.
@@ II Methods
Thin layers of hBN and graphite are exfoliated using adhesive tape on Si/SiO2 substrates. Subsequently, hBN/graphite heterostructures are assembled with a hot-pick-up technique [57] and cleaned by immersion in chloroform. The individual flake thickness is assessed with atomic force microscopy (AFM). To generate luminescent defects in hBN, we use a focused He-ion beam (ion energy of ) to pattern an array of circular patches with diameter and 2 µm pitch with ion fluences as defined in Table 1. Optical characterization is performed in a custom confocal microscope featuring a 532-nm laser and a grating spectrometer. Continuous-wave (cw) ODMR measurements with the PL integrated from 650 nm to 850 nm are performed with 532 nm laser excitation and a microwave excitation at 3 GHz – 4 GHz. Molecular dynamics (MD) simulations are carried out using the LAMMPS package [58], to extract the statistics of defects produced by the impacts of He ions onto the hBN. More details about the methods can be found in Appendix A and the motivation for the simulation setup is discussed in Refs. [59, 60, 61].
@@ III.1 Optical characterization of V–B centers
Figure 1 b depicts a typical hBN flake (80 nm thick) on graphite supported on Si/SiO2. The sample is exposed with arrays of circular patches of diameter and pitch. In those circular patches, the ion fluence is varied from to (see Table 1). To detect the defect emission, the PL is filtered from 650 nm to 850 nm and recorded with a single-photon counting device. The PL mapping after He-ion irradiation (Figure 1 c) clearly resolves the ion exposed spots by their increased luminescence (Figure A.2 a) as well as other bright spots, including folds and edges. A logarithmic scale is used for the color bar to visualize a larger dynamic range.
@@
Figure 1 d shows representative PL spectra recorded on the exposed spots for different ion fluences. The most prominent luminescence feature is a broad peak () centered around (), whose intensity increases monotonously with increasing He-ion exposure (Figure A.5 a). The PL can be saturated at large powers, about 5 mW for spots with (Figure A.5 b), which is indicative of emission from a finite number of localized defect centers [21, 28, 29, 30]. For comparison, the gray spectrum in Figure 1 d was recorded in a location not exposed to He-ions, i.e. on pristine hBN, where this spectral feature is absent. The peak position and spectral shape of the ion-induced luminescence feature are consistent with other studies that demonstrated emission from V–B [6, 17, 18, 19, 20, 21, 22, 23, 62].
@@
The large broadening, which obscures the observation of a zero-phonon line, is generally attributed to dominant phonon-assisted emission processes from an ensemble of emitters [6, 4, 63]. Cavity-enhanced experiments put the zero-phonon line of V–B at around 770 nm [48], whereas simulations show that phonon-independent luminescence is extremely weak [63]. It should be noted that making a quantitative comparison of the line shape for such a broad luminescence feature across different studies is often challenging, unless meticulous calibration of the overall spectral efficiency of the optics used is performed. This accounts for slight variations between this work and other works [6, 64, 28, 21, 30]. At the highest fluence (, dark purple spectrum in Figure 1 d), a background luminescence emerges across the full spectral range. This increased background correlates with the occurrence of an additional Raman mode at , which has previously been attributed to ion-beam induced damage in hBN [41, 28, 65] and which is absent in the pristine control spectrum (inset of Figure 1 d, see Figure A.1 the for full Raman spectrum). The background luminescence and defect-activated Raman mode may hint towards the onset of ion-induced amorphization of the hBN lattice [28] or the prevalent creation of larger defect clusters [65]. Consequently, up to the second highest fluence of , we can assume that luminescent defects are generated in a rather dilute limit.
@@
Consistently, we find that the PL lifetime does not display any clear change with irradiation fluence (or equivalently the induced defect density) except for the highest irradiation fluence, where the lifetime drops significantly from about to (Figure 2). The lifetimes are in line with values reported by other studies [6, 25, 31, 49, 30]. To fit the experimental data, a double-exponential model is applied, in which the first component captures the lifetime at longer time scales and the second component addresses the fine details at short time scales. The fit includes a convolution with the independently measured instrument response function. The time traces are shown in Figure 2 exemplarily for (red curve) and (blue curve). Table 2 in Appendix A.2 summarizes the extracted time constants (see Figure A.3 for the full lifetime data). Note that for optical emission with low quantum yield, as is the case for V–B, the measured lifetimes in time-resolved PL reflect a combination of the radiative and non-radiative rates. Therefore, the shortening of the lifetime with the fluence indicates an increased non-radiative recombination, which is likely due to the significant damage imparted onto the hBN lattice at the highest fluence. So far, studies reported tuning of the lifetime via temperature [31] or by exploiting spin-dependent non-radiative recombination using specific measurement pulse sequences [25, 30], where first long optical pulses initialize the spin polarization in the state and microwave -pulses rotate the spin polarization in one of the states ) thus shortening the overall lifetime. In our case we are therefore measuring the PL decay dynamics of a population in which both the radiative and non-radiative recombination mechanisms play a role [30].
@@ III.2 Optically detected magnetic resonance and microscopic charge model
The V–B defect is considered to be one of the few hBN defects whose atomic structure and level scheme are well known [6, 62, 64, 25, 31, 44, 52, 51]. Figure 3 e shows the level structure. The term indicates the projection of electron spin along the -axis of the hBN crystal, considered the quantization axis of the system [52]. The system exhibits a spin triplet ground state (GS) , with a so-called axial zero-field splitting around separating the and spin sublevels, and a corresponding spin triplet excited state (ES) , which is split between the spin sublevels as well [62, 64, 30]. Transitions between the ground and excited states are optically addressable under green laser excitation (typical wavelength 532 nm). Additionally, a manifold of metastable singlet states (MS) is present, which can be populated from the excited state via intersystem crossing (ISC). The phonon-assisted radiative emission from ES to GS is rather weak since it is forbidden by symmetry [30], thus these emitters display a low quantum efficiency. On the other side, recent studies found that the MS has a rather short lifetime (on the order of tens of nanoseconds) [29, 30]. Therefore, the defect can be pumped efficiently, and its PL can be detected at sufficient intensity despite the very low quantum efficiency [30].
@@
The ODMR spectra are measured at zero external magnetic field and on four different ion fluences (all on the same sample of Figure 1 c) corresponding to four emitter densities (Figure 3 a). All ODMR spectra display two characteristic resonances. We verified that their splitting increases in frequency under a static external magnetic field [6, 25] (Figure A.6). The slight difference in the overall magnitude of the contrast is due to a varying coupling efficiency of different positions in the sample to the nearby microwave antenna. To extract the splitting between the resonances, we fit the data with two exponentially modified Gaussians, which account for an asymmetric broadening of the resonances (see Figure A.7 b and the discussion in Appendix A.5). Figure 3 b shows the resonance positions and , extracted as the maxima of the individual resonance distribution, as a function of the irradiation fluence as well as their midpoint defined as .
@@
The splitting between the resonances increases with increasing ion fluence, or equivalently defect density (Figure 3 c). In literature, such splitting is attributed to the interaction of the V–B electronic spin with local fluctuations of the electric environment [44, 51, 52]. By contrast, variations of the mean are related to local strain fields [52] and temperature [66]. By modeling the magnitude of the splitting, it is possible to estimate the local field strength and, therefore, the local charge density surrounding the emitters. Following the approach in Refs. 44, 51, 52, we apply a microscopic charge model to simulate the level splitting. Briefly, we consider level splitting due to a net electric field, which is generated by a discrete and random charge distribution in the vicinity of the V–B site. Importantly, the model still assumes overall charge neutrality, i.e. positive and negative point charges are placed around the V–B in equal numbers (see Appendix A.5 for details). The remaining net electric field arises solely from the locally net sum of the field components. To date, the microscopic origin of such a charge environment is not fully understood. A common assumption is that the charge fluctuations simply correspond to the number of charged V–B defects, which, as we discuss in the following, is not fully consistent with our experimental results.
@@
Generally, the extracted charge densities are in line with other reports [44, 51, 52]. This finding is consistent with the fact that we measure similar magnitudes of the splitting and incorporate the value of the out-of-plane susceptibility determined in Ref. 44 as the main parameter in the model. However, unlike Ref. 52, we do not observe a proportional scaling of charge density and PL intensity with ion fluence, whereby the latter can be taken as a good estimate of the relative number of optically active defects for a constant lifetime (cf. Figure 2) and for an excitation power below the saturation threshold (Figure A.5). Specifically, the PL intensity of V–B emission increases by a factor of 10 (Figure A.5) as the ion fluence is varied over more than one order of magnitude from to while, at the same time, the average charge density extracted from the model increases only slightly from to . Therefore, as a possible extension of the model, we propose that the overall charge density should include a constant positive and negative background charge, which is present before ion irradiation. The origin of such a background charge may potentially be related to native defects, such as interstitials, impurities, or vacancies present after the growth. Charge transfer processes between such defects can then lead to local charge accumulation, while preserving overall charge neutrality. Such a background charge would lead to a finite level splitting already in the limit of small V–B defect density, as we clearly observe for (Figure 3 a). With these assumptions, we determine a background charge of . From that, we can find the corrected values for the V–B density as given in Table 1.

### v3

# Quantifying the creation of negatively charged boron vacancies in He-ion irradiated hexagonal boron nitride

@@ Abstract
Hexagonal boron nitride (hBN) hosts luminescent defects possessing spin qualities compatible with quantum sensing protocols at room temperature. Vacancies, in particular, are readily obtained via exposure to high-energy ion beams. While the defect creation mechanism via such irradiation is well understood, the occurrence rate of optically active negatively charged vacancies (V–B) is still an open question. In this work, we exploit focused helium ions to systematically create optically active vacancy defects in hBN flakes at varying density. By comparing the density-dependent spin splitting measured by magnetic resonance to calculations based on a microscopic charge model, in which we introduce a correction term due to a constant background charge, we are able to quantify the number of V–B defects created by the ion irradiation. We find a lower bound for the fraction (0.2%) of all vacancies in the optically active, negatively charged state. Our results provide a protocol for measuring the creation efficiency of V–B, which is necessary for understanding and optimizing luminescent centers in hBN.
@@ I Introduction
Luminescent centers in hBN emerged as bright and stable room temperature quantum emitters [1, 2], rivaling established platforms in quantum communication and sensing applications [3, 4, 5]. Emitters in hBN exhibit robust photoluminescence (PL) across a broad spectral range from the near-infrared [6] through the visible [1, 7] up to the blue [8, 9, 10, 11, 12] and UV [13] regions, as well as magnetic response both as ensemble spin-1 systems [6] and as single spin-1 [14, 15] or spin-1/2 [16] emitters. Among the most studied types of luminescent centers in hBN is the negatively charged boron vacancy (V–B), which consists of an electron localized at the position of the missing atom in the hexagonal lattice. These defects are typically addressed in PL ensemble measurements, where they exhibit a very broad spectrum centered around 800 nm [6, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Importantly, V–B defects feature a spin-triplet ground state with a zero-field splitting of around 3.5 GHz [6], a corresponding triplet excited state with a splitting of around 2.1 GHz [25, 31, 32, 33], and a metastable state acting as a non-radiative channel with a very short lifetime [29, 30]. The spin state can be optically read-out and coherently manipulated through optically detected magnetic resonance (ODMR) measurement schemes [6, 28, 34, 35, 30]. The spin sub-levels, and thus the two resonances in the ODMR spectrum, are sensitive to external magnetic fields [6, 25], which can be applied for quantum sensing schemes [36, 37, 38, 18], similar to nitrogen vacancy-centers in diamond [39, 40]. Because of their potential for applications in quantum technologies, the optical and magnetic properties of V–B defects have been a subject of intensive research, including their optical response at room [6, 25, 41, 28, 30] and cryogenic [20, 41] temperatures, as well as ODMR [6, 20, 31, 33], electron paramagnetic resonance (EPR) [6, 42], and spin-coherence measurements [4, 43, 44, 45]. Furthermore, several studies discuss various sample preparation and optimization protocols, such as irradiation sputtering [22], ion beam fluence and energy tuning [21, 41], high-temperature annealing [46, 23, 24], water exposure [22], as well as the integration of defects into photonic circuits, such as nanocavities [47, 48], plasmonic- [19] and metasurfaces [49].
@@
For photonic integration reliable, quantitative and deterministic methods for defect creation are desirable. In this context, different methods were successfully applied to create V–B in hBN, which include focused ion beam irradiation using H [17], He [46, 28, 41, 19, 23, 21, 50], N [30, 21, 20, 51, 52], Ga [6, 31], C [21, 25, 53], Xe [20] and Ar [21, 20], neutron irradiation [27, 6, 38, 54, 55, 56], femtosecond laser writing [57], and electron beam irradiation [42, 58]. In particular, several works discussed the usage of a focused He-ion beam source to effectively create V–B defects in hBN [28, 41, 19, 23, 21, 50]. Focused He-ion beam exposure offers a nominal lateral defect positioning accuracy below 10 nm in very thin films [59], while preserving the local crystal structure of hBN at adequate doses. All such nanofabrication measures are crucial factors for an efficient integration of such luminescent centers in future photonic [60] and quantum imaging devices [18]. In this context, several works [41, 28, 23] already provided initial insights onto the He-ion creation process and the optical response of the created ensemble emitters, but a systematic understanding, in particular a reliable quantification of defect density and its impact on the magnetic and optical properties of the created emitters, is still lacking.
@@
In the present work, we use a helium ion microscope (HIM) to irradiate hBN flakes, deposited on thick graphite flakes, to create V–B defects (Figure 1 a). The thick graphite acts as both a mechanical and optical protection layer, reducing the impact of back-scattered ions from the SiO2 substrate and suppressing unwanted background luminescence from the irradiation damage in the substrate. We systematically perform spatially- and time-resolved PL as well as ODMR measurements on defect ensembles for varying irradiation fluences, confirming the optical and magnetic activity of the created emitters. We show that the focused He-ion beam technique can site-selectively produce luminescent and background-free emitters, in good agreement with the existing literature [6, 28, 41, 19, 23, 21, 50]. Furthermore, we apply a microscopic charge model [44, 55] to describe the defect density-dependent shift of the ODMR resonance frequencies. By introducing a background charge correction, we can quantify from the experimental data the density of the charged defects created by the ion exposure. By comparison to molecular dynamics (MD) simulations, which estimate the number of total vacancies created, we are able to quantify an overall efficiency of the charged defect creation process based on the focused He-ion beam irradiation. Our work provides a systematic study of defect production efficiency by comparing different irradiation fluences on homogeneous and uniform samples, offering a benchmark for future process optimization and device implementation.
@@ II Methods
Thin layers of hBN and graphite are exfoliated using adhesive tape on Si/SiO2 substrates. Subsequently, hBN/graphite heterostructures are assembled with a hot-pick-up technique [61] and cleaned by immersion in chloroform. The individual flake thickness is assessed with atomic force microscopy (AFM). To create luminescent defects in hBN, we use a focused He-ion beam (ion energy of ) to pattern an array of circular patches with diameter and 2 µm pitch with ion fluences as defined in Table 1. Optical characterization is performed in a custom confocal microscope featuring a 532-nm laser and a grating spectrometer. Continuous-wave (cw) ODMR measurements, with the PL integrated from 650 nm to 850 nm, are performed with 532 nm laser excitation and a microwave excitation at 3 GHz – 4 GHz. Molecular dynamics (MD) simulations are carried out using the LAMMPS package [62], to extract the statistics of defects produced by the impacts of He ions onto the hBN. More details about the methods can be found in Appendix A and the motivation for the simulation setup is discussed in Refs. [63, 64, 65].
@@ III.1 Optical characterization of V–B centers
Figure 1 b depicts a typical hBN flake (80 nm thick) on graphite supported on Si/SiO2. The sample is exposed with arrays of circular patches of diameter and pitch. In those circular patches, the ion fluence is varied from to (see Table 1). To spatially map the defect emission, we use confocal fluorescence imaging. Here, the PL is filtered from 650 nm to 850 nm and recorded with an avalanche photodiode in photon counting mode while the excitation laser is scanned across the sample. The fluorescence mapping after He-ion irradiation (Figure 1 c) clearly resolves the ion-exposed spots by their increased luminescence (Figure A.2 a) as well as other bright spots, including folds and edges. A logarithmic scale is used for the color bar to visualize a larger dynamic range.
@@
Figure 1 d shows representative PL spectra recorded on the exposed spots for different ion fluences. The most prominent luminescence feature is a broad peak () centered around (), whose intensity increases monotonously with increasing He-ion exposure (Figure A.5 a). Consistent with earlier reports [20, 19, 30], the PL can be saturated at large powers, about 1.37 mW (measured at the entrance aperture of the objective) for spots with (Figure A.5 b), which is indicative of emission from a finite number of localized defect centers [21, 28, 29, 30]. For comparison, the gray spectrum in Figure 1 d was recorded in a location not exposed to He-ions, i.e. on pristine hBN, where this spectral feature is absent. The peak position and spectral shape of the ion-induced luminescence feature are consistent with other studies that demonstrated emission from V–B [6, 17, 18, 19, 20, 21, 22, 23, 66].
@@
The large broadening, which obscures the observation of a zero-phonon line, is generally attributed to dominant phonon-assisted emission processes from an ensemble of emitters [6, 4, 67]. Cavity-enhanced experiments put the zero-phonon line of V–B at around 770 nm [48], whereas simulations show that phonon-independent luminescence is extremely weak [67]. It should be noted that making a quantitative comparison of the line shape for such a broad luminescence feature across different studies is often challenging, unless meticulous calibration of the overall spectral efficiency of the optics used is performed. This accounts for slight variations between this work and other works [6, 68, 28, 21, 30]. At the highest fluence (, dark purple spectrum in Figure 1 d), a background luminescence emerges across the full spectral range. This increased background correlates with the occurrence of an additional Raman mode at , which has previously been attributed to ion-beam induced damage in hBN [41, 28, 69] and which is absent in the pristine control spectrum (inset of Figure 1 d, see Figure A.1 for the full Raman spectrum). The background luminescence and defect-activated Raman mode may hint towards the onset of ion-induced amorphization of the hBN lattice [28] or the prevalent creation of larger defect clusters [69].
@@
Consistently, we find that the PL lifetime does not display any clear change with irradiation fluence (or equivalently the induced defect density) except for the highest irradiation fluence, where the lifetime drops significantly from about to (Figure 2). The lifetimes are in line with values reported by other studies [6, 25, 31, 49, 30]. To fit the experimental data, a double-exponential model is applied, in which the first component captures the lifetime at longer time scales and the second component addresses the fine details at short time scales. The fit includes a convolution with the independently measured instrument response function. The time traces are shown in Figure 2 exemplarily for (red curve) and (blue curve). Table 2 in Appendix A.2 summarizes the extracted time constants (see Figure A.3 for the full lifetime data). Note that for optical emission with low quantum yield, as is the case for V–B, the measured lifetimes in time-resolved PL reflect a combination of the radiative and non-radiative rates. Therefore, the shortening of the lifetime with the fluence indicates an increased non-radiative recombination, which is likely due to the significant damage imparted onto the hBN lattice at the highest fluence. So far, studies reported tuning of the lifetime via temperature [31] or by exploiting spin-dependent non-radiative recombination using specific measurement pulse sequences [25, 30] (where first long optical pulses initialize the spin polarization in the state and microwave -pulses rotate the spin polarization in one of the states ), thus shortening the overall lifetime. In our case we are therefore measuring the PL decay dynamics of a population in which both the radiative and non-radiative recombination mechanisms play a role [30].
@@ III.2 Optically detected magnetic resonance and microscopic charge model
The V–B defect is considered to be one of the few hBN defects whose atomic structure and level scheme are well known [6, 66, 68, 25, 31, 44, 55, 54]. Figure 3 e shows the level structure. The system exhibits a spin triplet ground state (GS) and a corresponding spin triplet excited state (ES) . The ground state shows a so-called axial zero-field splitting around separating the and spin sublevels. The term indicates the projection of electron spin along the -axis of the hBN crystal, considered the quantization axis of the system [55]. An electric field perpendicular to the quantization axis mixes the sub-levels, resulting in new eigenstates and with a level splitting of (Figure 3 e inset), where denotes the corresponding perpendicular susceptibility. Transitions between the ground and excited states are optically addressable under green laser excitation (typical wavelength 532 nm). Additionally, a manifold of metastable singlet states (MS) is present, which can be populated from the excited state via intersystem crossing (ISC). The phonon-assisted radiative emission from excited state to ground state is rather weak since it is forbidden by symmetry [30], thus V–B emitters intrinsically have low radiative recombination rate and, hence, a low quantum efficiency. At the same time, the metastable state has a rather short lifetime (on the order of tens of nanoseconds) [29, 30]. Therefore, the systems relaxes overall rather rapidly, and the defect can be re-pumped efficiently, such that its PL can still be detected at sufficient intensity despite the intrinsically low quantum efficiency and the presence of the metastable state [30].
@@
The ODMR spectra are measured at zero external magnetic field and on four different ion fluences (all on the same sample of Figure 1 c) corresponding to four emitter densities (Figure 3 a). All ODMR spectra display two characteristic resonances. We verified that their splitting increases in frequency under a static external magnetic field [6, 25] (Figure A.6). The slight difference in the overall magnitude of the contrast is due to a varying coupling efficiency of different positions in the sample to the nearby microwave antenna. To extract the splitting between the resonances, we fit the data with two exponentially modified Gaussians, which account for an asymmetric broadening of the resonances (see Figure A.7 b and the discussion in Appendix A.5). Figure 3 b shows the resonance positions and , extracted as the maxima of the individual resonance distribution, as a function of the irradiation fluence, as well as their midpoint defined as .
@@
The splitting between the resonances increases with increasing ion fluence, or equivalently defect density (Figure 3 c). In literature, such splitting is attributed to the interaction of the V–B electronic spin with local fluctuations of the electric environment [44, 54, 55]. By contrast, variations of the mean are related to local strain fields [55] and temperature [70]. By modeling the magnitude of the splitting, it is possible to estimate the local field strength and, therefore, the local charge density surrounding the emitters. Following the approach in Refs. 44, 54, 55, we apply a microscopic charge model to simulate the level splitting. Briefly, we consider level splitting due to a net electric field, which is generated by a discrete and random charge distribution in the vicinity of the V–B site. Importantly, the model still assumes overall charge neutrality, i.e. positive and negative point charges are placed around the V–B in equal numbers (see Appendix A.5 for details). The remaining net electric field arises solely from the locally net sum of the field components. To date, the microscopic origin of such a charge environment is not fully understood. A common assumption is that the charge fluctuations simply correspond to the number of charged V–B defects, which, as we discuss in the following, is not fully consistent with our experimental results.
@@
Generally, the extracted charge densities are in line with other reports [44, 54, 55]. This finding is consistent with the fact that we measure similar magnitudes of the splitting and incorporate the value of the out-of-plane susceptibility determined in Ref. 44 as the main parameter in the model. However, unlike Ref. 55, we do not observe a proportional scaling of charge density and PL intensity with ion fluence, whereby the latter can be taken as a good estimate of the relative number of optically active defects for a constant lifetime (cf. Figure 2) and for an excitation power below the saturation threshold (Figure A.5). Specifically, the PL intensity of V–B emission increases by a factor of 10 (Figure A.5) as the ion fluence is varied over more than one order of magnitude from to while, at the same time, the average charge density extracted from the model increases only slightly from to . Therefore, as a possible extension of the model, we propose that the overall charge density should include a constant positive and negative background charge, potentially due to other defect species. Charge transfer processes between such defects can then lead to local charge accumulation, while preserving overall charge neutrality. Such a background charge would lead to a finite level splitting already in the limit of small V–B defect density, as we clearly observe for (Figure 3 a). With these assumptions, we determine a background charge of . From that, we can find the corrected values for the V–B density as given in Table 1. The origin of such a background charge may potentially be related to native defects, such as interstitials, impurities, or vacancies present after the crystal growth and, hence, before the irradiation process. Additionally, defects introduced already at very low irradiation fluences, below the rather large fluences needed to create enough optically active V–B sites for reliable ODMR characterization, may contribute substantially to the observed ODMR splitting (see below for a more detailed discussion on the impact of crystal quality).